# Will the Real Linda Please Stand up…to Large Language Models? Examining the Representativeness Heuristic in LLMs

### v1

@@ Abstract
Although large language models (LLMs) have demonstrated remarkable proficiency in understanding text and generating human-like text, they may exhibit biases acquired from training data in doing so. Specifically, LLMs may be susceptible to a common cognitive trap in human decision-making called the representativeness heuristic. This is a concept in psychology that refers to judging the likelihood of an event based on how closely it resembles a well-known prototype or typical example versus considering broader facts or statistical evidence. This work investigates the impact of the representativeness heuristic on LLM reasoning. We created ReHeAT (Representativeness Heuristic AI Testing), a dataset containing a series of problems spanning six common types of representativeness heuristics. Experiments reveal that four LLMs applied to ReHeAT all exhibited representativeness heuristic biases. We further identify that the model’s reasoning steps are often incorrectly based on a stereotype rather than the problem’s description. Interestingly, the performance improves when adding a hint in the prompt to remind the model of using its knowledge. This suggests the uniqueness of the representativeness heuristic compared to traditional biases. It can occur even when LLMs possess the correct knowledge while failing in a cognitive trap. This highlights the importance of future research focusing on the representativeness heuristic in model reasoning and decision-making and on developing solutions to address it.
@@ 1 Introduction
This is one experiment known as the “Linda problem” devised by Tversky and Kahneman; it was demonstrated that people are influenced by specific descriptions, such as being deeply concerned with issues of discrimination and social justice and also participating in anti-nuclear demonstrations to rank [7] higher than [5]. However, [7] combines [5] and an additional event. This means that, from a statistical perspective, [5] is more likely to occur than [7] because it is more general and less restrictive.
@@
Kahneman & Tversky (1973) introduced this phenomenon as the “representativeness heuristic,” which involves estimating the likelihood of an event by comparing it to an existing prototype in our minds. It offers a convenient shortcut in decision-making by aligning with intuitive thinking, leading people to rely on it frequently. Generally, this heuristic is quite beneficial, as it simplifies complex judgments. However, it is important to recognize that it can also result in significant errors, given that people are prone to assessing the likelihood of an object’s category membership based on superficial similarities while neglecting actual statistical evidence. For example, people categorize individuals based on their looks, actions, or background description from language, leading to skewed perceptions and decisions. This stereotyping phenomenon is widespread among humans (Spencer et al., 2016).
@@
LLMs trained on real-world data and instructed to emulate human behavior may capture the representativeness heuristic. Previous work has mainly focused on the biases within training data (Bender et al., 2021; Garg et al., 2018a; Sheng et al., 2019; Zhao et al., 2019; Bolukbasi et al., 2016a). These biases often stem from data distributions that do not reflect the desired proportions that would drive unbiased decision-making. In contrast, the representativeness heuristic represents a type of cognitive bias that has yet been thoroughly investigated in LLMs. It is unique in leading the model to make mistakes even when it possesses the knowledge necessary to solve the problem. As illustrated in Figure 1, the model is able to answer the Statistical Prototype Question (SPQ), yet it tends to fail to answer the Representativeness Heuristic Question (RHQ). The SPQ and RHQ are intrinsically equivalent, where SPQ is expressed statistically, and RHQ expresses the same statistical logic in a scenario. This indicates that LLMs can engage in erroneous cognitive reasoning even with the knowledge of statistical probability. Interestingly, providing a hint can prompt the model to use the knowledge to make a correct prediction. This indicates that the representativeness heuristic can block the model from following a correct reasoning path; instead, it relies on the cognitive shortcut to make a trivial decision.
@@
Figure 1: Illustration of the representativeness heuristic problem. The model possesses the knowledge to answer the statistical prototype question but fails to use it to solve the representativeness heuristic question. Providing a hint in the prompt can guide the model to make a correct prediction.
@@
To investigate the representativeness heuristic in LLMs, we construct a dataset [^1] We will release the dataset upon the completion of the review process to facilitate future research. , ReHeAT (Representativeness Heuristic AI Testing), which contains 202 RHQs that span six types of representativeness heuristics—Base Rate Fallacy, Conjunction Fallacy, Disjunction Fallacy, Insensitivity to Sample Size, Misconceptions of Chance, and Regression Fallacy. The questions we designed are adapted from those used in prior investigations into heuristics within the field of psychology (Bar-Hillel & Neter, 1993; Kahneman & Tversky, 1973; Kahneman et al., 1982; Tversky & Kahneman, 1974; 1983). To the best of our knowledge, our dataset is the first to offer extensive and comprehensive coverage of RHQs, enabling exploration of LLMs’ capabilities in countering the cognitive bias.
@@
We evaluate four LLMs: GPT-3.5 (Ouyang et al., 2022), GPT-4 (OpenAI, 2023), PaLM 2 (Anil et al., 2023), and LLaMA 2 (Touvron et al., 2023), on ReHeAT using different prompts. Our findings indicate that these LLMs exhibit behaviors that closely mirror human heuristic behavior. Additionally, advanced prompting techniques such as chain of thought (CoT) (Wei et al., 2022), in-context learning (Brown et al., 2020), and self-consistency (Huang et al., 2022) prompting, offer marginal improvements. Nevertheless, when explicitly prompted to recall its knowledge, the model shows an improvement in performance. This underscores the significance for future research to address the representativeness heuristic, guiding LLMs toward correct reasoning and decision-making.
@@ 2 Related Work
Social biases in natural language processing (NLP) systems and related data have been studied with respect to their fairness, inclusivity, and accuracy  (Maass, 1999; Zhao et al., 2018; Hutchinson et al., 2020). For example, Bolukbasi et al. (2016b); Garg et al. (2018b) are among the pioneers in demonstrating gender biases embedded in word embeddings, showing how these representations could reinforce stereotypical associations. Caliskan et al. (2017) concludes standard machine learning methods for NLP could acquire societal biases from textual data. Some works expanded the understanding of bias sources in NLP systems, including those from data collection (Bender & Friedman, 2018), annotation processes (Gebru et al., 2018), and model architecture choices (Zhao et al., 2017). To date, numerous efforts have been made to mitigate social biases in systems through a variety of methods, including data augmentation (Lu et al., 2018), changes in model architecture (Liang et al., 2020), and training objectives (Romanov et al., 2019; Liu et al., 2021).
In a similar vein, although recent advancements in LLMs are exciting, researchers are concerned about whether LLMs inherit social biases from the trillions of tokens they have been trained on. Weidinger et al. (2022) gives a comprehensive taxonomy of social risks within LLMs. Although the research community has documented numerous social biases in LLMs (Mei et al., 2023; Ferrara, 2023), few LLM researchers have examined these biases from the human mind’s perspective. In this work, we study the bias issue in LLMs from a new angle, representative heuristics, a concept originating in psychology (Kahneman & Tversky, 1973; Kahneman et al., 1982; Tversky & Kahneman, 1974; 1983).
@@ 3 Representativeness Heuristic
Misconceptions of Chance involve misunderstanding how randomness works, especially thinking that past outcomes will affect future outcomes, in cases where the outcomes are in fact independent. For example, Figure 2(e) presents a dice-rolling example. People wrongly believe that if a specific outcome has occurred frequently, it is less likely to happen again shortly, or vice versa if it has occurred rarely. However, in truly random events, such as rolling a fair die, the probability of any given outcome (1-6) remains constant at $\frac{1}{6}$, unaffected by the sequence of previous results. An example question can be seen in Table 21 of the Appendix.
@@ 4.1 Data Collection
Our main resource for creating test questions on representativeness heuristic comes from the academic studies by Kahneman and Tversky (Kahneman & Tversky, 1973; Kahneman et al., 1982; Tversky & Kahneman, 1974; 1983). Their work introduced various question types and insights into the design of cognitive heuristic measures. Building on their foundation, we design questions that extensively explore the representativeness heuristic in LLMs. Specifically, our test set contains 49 questions drawn directly from previous research and 153 new questions that have been carefully adapted. To support the usefulness of our adapted items, we pay special attention to retaining the essence of the representativeness heuristic being tested, while changing the situation and context of the original items.
@@
    Type: Indicate the question’s category, such as Disjunction Fallacy.
@@
    Feature: Highlight unique aspects, such as choose one, or the question’s source, such as original problem in Bar-Hillel & Neter, 1993.
@@
Feature Choose one, original problem in Bar-Hillel & Neter, 1993
@@
Table 1: Examplar in ReHeAT
@@ 5.1 LLMs Exhibit the Representativeness Heuristic
We also check each type of representativeness heuristic (Table 6 - 9 of the Appendix). We have observed that in most cases, most models perform relatively poorly when dealing with questions of Conjunction Fallacy and Disjunction Fallacy. Compared to other types of questions, such as Base Rate Fallacy, the challenge with Conjunction Fallacy and Disjunction Fallacy lies in the fact that the statistical reasoning required is embedded within the connotations and combinations of texts (e.g., South Africa is a subset of Africa) rather than being directly indicated by expressions like large sample size or far more than.
@@
Overall, LLMs have demonstrated a representativeness heuristic bias similar to humans, as found in the research findings reported by Bar-Hillel & Neter (1993); Kahneman & Tversky (1973); Kahneman et al. (1982); Tversky & Kahneman (1974; 1983), where human accuracy across various tasks was within the 10% to 50% range. Example questions and model answers are in Table 17 - 22 of the Appendix.
@@ 5.2 Discrepancies in Model Reasoning: Beyond Predictions
Generally, GPT-4 outperformed LLaMA2-70B in all types of RHQs. However, this improvement is accompanied by a side effect: an increased proportion of instances where GPT-4’s reasoning process is accurate, but the outcome is incorrect. This primarily occurs because when faced with ambiguous questions, GPT-4 often preferred to express the need for additional information before making a decision. This is also why GPT-4 accuracy is slightly lower than GPT-3.5 under some prompts. Although we mark such cautious answers as incorrect outcomes—because they did not adhere to the directive of providing a definitive answer—cautious answers may remain beneficial in practical settings. Conversely, LLaMA2-70B more frequently produces correct outcomes from incorrect reasoning processes; the reasons are often incorrectly based on a stereotype rather than the problem’s description. We have conducted case studies on some interesting reasoning, we include the case study in Appendix A.
@@
5.3 Improving Performance by Hinting that LLMs to use Their Knowledge
How can the model’s cognitive process be put back on track? We test whether LLMs show enhanced performance with prompts that hint them to utilize their existing knowledge. We tested two types of hints: one general and the other based on more detailed cues for each representativeness heuristic type (Table 16 of the Appendix). These prompts aimed to hint at the model to recall the knowledge it possessed. The results are presented in Table 3. We also provide detailed results for each type of representativeness heuristic in Table 11 and Table 12 of the Appendix. We found that both types of hints provided a noticeable improvement for most models, with the specific type of hints providing a more significant boost in performance compared to the general prompts.
@@ 5.4 The Impact of Situational Similarity on Model Performance
After observing the performance differences between the model on statistical SPQs and the contextual RHQ counterparts found in ReHeAT, we delve further into the transition between these two - Intermediate Questions (IQs) (Table 15 of the Appendix). This type of question integrates specific situational contexts and statistical data, making it necessary to consider both concrete data and the potential impact of the situation when making decisions. For example, for Linda’s problem, we assign a probability to each independent event to offer the model with more explicit statistical information. Nevertheless, the model has to infer the information based on the similarity manifested in the meaning of the statements in RHQs. The performance of the models on IQs is reported in Table 5 of the Appendix. Compared with SPQs’ results (Table 4 of the Appendix), more errors were made for IQs. This indicates that introducing scenarios interferes with the model’s statistical decision-making process. This also explains why LLMs exhibit the representativeness heuristic more frequently when responding to RHQs.

### v4

@@ Abstract
Although large language models (LLMs) have demonstrated remarkable proficiency in modeling text and generating human-like text, they may exhibit biases acquired from training data in doing so. Specifically, LLMs may be susceptible to a common cognitive trap in human decision-making called the representativeness heuristic. This is a concept in psychology that refers to judging the likelihood of an event based on how closely it resembles a well-known prototype or typical example, versus considering broader facts or statistical evidence. This research investigates the impact of the representativeness heuristic on LLM reasoning. We created ReHeAT (Representativeness Heuristic AI Testing), a dataset containing a series of problems spanning six common types of representativeness heuristics. Experiments reveal that four LLMs applied to ReHeAT all exhibited representativeness heuristic biases. We further identify that the model’s reasoning steps are often incorrectly based on a stereotype rather than on the problem’s description. Interestingly, the performance improves when adding a hint in the prompt to remind the model to use its knowledge. This suggests the uniqueness of the representativeness heuristic compared to traditional biases. It can occur even when LLMs possess the correct knowledge while falling into a cognitive trap. This highlights the importance of future research focusing on the representativeness heuristic in model reasoning and decision-making and on developing solutions to address it.
@@ 1 Introduction
This is one experiment known as the “Linda problem,” devised by Tversky and Kahneman; results demonstrate how respondents are influenced by specific descriptions, such as being deeply concerned with issues of discrimination and social justice and also participating in anti-nuclear demonstrations to rank [7] higher than [5]. However, [7] combines [5] and an additional event. This means that, from a statistical perspective, [5] is more likely to occur than [7] because it is more general and less restrictive.
@@
Kahneman & Tversky (1973) introduced this phenomenon as the “representativeness heuristic,” which involves estimating the likelihood of an event by comparing it to an existing prototype in our minds. It offers a convenient shortcut in decision-making by aligning with intuitive thinking, leading people to rely on it frequently. Generally, this heuristic is quite beneficial, as it simplifies complex judgments. However, it is important to recognize that it can also result in significant decision-making errors, given that people are prone to assessing the likelihood of an object’s category membership based on superficial similarities, while neglecting actual statistical evidence. For example, when they lack the benefit of individuating information about a person, decision-makers might categorize individuals based on their physical appearance, observation of limited behaviors, or superficial background descriptions, leading to skewed perceptions and decisions. This type of stereotyping phenomenon is durable, being widespread across cultures and over time (Spencer et al., 2016).
@@
Because LLMs are trained on real-world data and instructed to emulate human behavior, they may capture the representativeness heuristic. Previous work has mainly focused on the biases within training data (Bender et al., 2021; Bolukbasi et al., 2016; Garg et al., 2018; Sheng et al., 2019; Zhao et al., 2019). These biases often stem from data distributions that do not reflect proportions that would drive decision-making that is unbiased, given a researcher’s operational definition of bias. In contrast, the representativeness heuristic represents a type of cognitive bias that has yet to be thoroughly investigated in LLMs. It is unique in potentially leading models to make mistakes, even when they possess the knowledge necessary to solve the problem. As illustrated in Figure 1, the model is able to answer the Statistical Prototype Question (SPQ), yet it tends to fail to answer the Representativeness Heuristic Question (RHQ). The SPQ and RHQ are intrinsically equivalent, where SPQ is expressed statistically, and RHQ expresses the same statistical logic verbally in a scenario. This indicates that LLMs can engage in erroneous cognitive reasoning, even with accurate knowledge of statistical probabilities. Interestingly, providing an appropriate hint can prompt an LLM to use its knowledge to make correct predictions more frequently. This indicates that the representativeness heuristic can block the model from following a correct reasoning path; instead, it relies on the data that captures the human tendency to use cognitive shortcuts to make decisions.
@@
Figure 1: Illustration of the representativeness heuristic problem. The model possesses the knowledge to answer the statistical prototype question, yet fails to use it to solve the representativeness heuristic question. Providing an appropriate hint in the prompt can guide the model in making a correct prediction.
@@
To investigate the representativeness heuristic in LLMs, we construct a dataset [^1] Code and dataset will be available at https://github.com/MrZilinXiao/LLMHeuristicReHEAT. , ReHeAT (Representativeness Heuristic AI Testing), which contains 202 RHQs that span six types of representativeness heuristics: Base Rate Fallacy, Conjunction Fallacy, Disjunction Fallacy, Insensitivity to Sample Size, Misconceptions of Chance, and Regression Fallacy. The questions we designed are adapted from those used in prior investigations into heuristics within the field of psychology (Bar-Hillel & Neter, 1993; Kahneman & Tversky, 1973; Kahneman et al., 1982; Tversky & Kahneman, 1974; 1983). To the best of our knowledge, our dataset is the first to offer extensive and comprehensive coverage of RHQs, enabling exploration of LLMs’ capabilities in countering the representative heuristic cognitive bias.
@@
We evaluate four LLMs: GPT-3.5 (Ouyang et al., 2022), GPT-4 (OpenAI, 2023), PaLM 2 (Anil et al., 2023), and LLaMA 2 (Touvron et al., 2023), on ReHeAT using different prompts. Our findings indicate that these LLMs exhibit behaviors that closely mirror human heuristic behavior. Additionally, advanced prompting techniques such as chain of thought (CoT) (Wei et al., 2022), in-context learning (Brown et al., 2020), and self-consistency (Huang et al., 2022) prompting, offer marginal improvements. Nevertheless, when explicitly prompted to recall its knowledge, the model shows an improvement in performance. This underscores the significance of future research to address the representativeness heuristic, guiding LLMs toward correct reasoning and decision-making.
@@ 2 Related Work
Social biases in natural language processing (NLP) systems and related data have been studied with respect to their fairness, inclusivity, and accuracy  (Hutchinson et al., 2020; Maass, 1999; Zhao et al., 2018). For example, Bolukbasi et al. (2016); Garg et al. (2018) are among the pioneers in demonstrating gender-related associations in word embeddings that might reflect and perpetuate stereotypes. Caliskan et al. (2017) concludes standard machine learning methods for NLP could acquire societal biases from textual data. Some research has expanded our understanding of where NLP systems might acquire subgroup associations within data, and potential bias, including those from data collection (Bender & Friedman, 2018), annotation processes (Gebru et al., 2018), and model architecture choices (Zhao et al., 2017). To date, numerous efforts have been made to mitigate social biases in systems through a variety of methods, including data augmentation (Lu et al., 2018), changes in model architecture (Liang et al., 2020), and training objectives (Liu et al., 2021; Romanov et al., 2019).
In a similar vein, although recent advancements in LLMs are exciting, researchers are concerned about whether LLMs inherit social biases from the trillions of tokens they have been trained on. Weidinger et al. (2022) provides a comprehensive taxonomy of social risks within LLMs. Although the research community has documented numerous social biases in LLMs (Ferrara, 2023; Mei et al., 2023), few LLM researchers have examined these biases from the perspective of the psychology of human decision-making. Thus, in the present work, we study the bias issue in LLMs from the new angle of the aforementioned representative heuristic, a concept originating in psychology (Kahneman & Tversky, 1973; Kahneman et al., 1982; Tversky & Kahneman, 1974; 1983).
@@ 3 Representativeness Heuristic
Misconceptions of Chance involve misunderstanding how randomness works, especially thinking that past outcomes will affect future outcomes in cases where the outcomes are, in fact, independent. For example, Figure 2(e) presents a dice-rolling example. People wrongly believe that if a specific outcome has occurred frequently, it is less likely to happen again shortly, or vice versa if it has occurred rarely. However, in truly random events, such as rolling a fair die, the probability of any given outcome (1-6) remains constant at $\frac{1}{6}$, unaffected by the sequence of previous results. An example question can be seen in Table 22 of the Appendix.
@@ 4.1 Data Collection
Our main resource for creating test questions on representativeness heuristic comes from academic research by Kahneman and Tversky (Kahneman & Tversky, 1973; Kahneman et al., 1982; Tversky & Kahneman, 1974; 1983). Their work introduced various question types and insights into the design of cognitive heuristic measures. Building on their foundation, we design questions that extensively explore the representativeness heuristic in LLMs. Specifically, our test set contains 49 questions drawn directly from previous research and 153 new questions that have been carefully adapted. The reason for creating these new questions is to provide diverse contexts and achieve stronger external validity. In creating our adapted items that change the situation and context of the original items, we paid special attention to retaining the essence of the representativeness heuristic being tested.
@@
    Type: Indicate the question’s category, such as: Disjunction Fallacy.
@@
    Feature: Highlight unique aspects, such as choose one, or the question’s source, such as: original problem in Bar-Hillel & Neter (1993).
@@
Feature Choose one, original problem in Bar-Hillel & Neter (1993)
@@
Table 1: Exemplar in ReHeAT
@@ 5.1 LLMs Exhibit the Representativeness Heuristic
We also check each type of representativeness heuristic (Table 7 - 10 of the Appendix). We have observed that in most cases, most models perform relatively poorly when dealing with questions of Conjunction Fallacy and Disjunction Fallacy. Compared to other types of questions, such as Base Rate Fallacy, the challenge with Conjunction Fallacy and Disjunction Fallacy lies in the fact that the statistical reasoning required is embedded within the connotations and combinations of texts (e.g., South Africa is a subset of Africa) rather than being directly indicated by expressions like large sample size or far more than.
@@
Human Performance on ReHeAT
We also present human performance on ReHeAT in Table 3. Through the Prolific platform, we recruited 153 participants (66 males, 86 females, and 1 other) to answer questions. The participants’ ages ranged from 19 to 77 years (M = 37.2, SD = 12.6). The racial distribution of the participants was as follows: 70.6% White, 9.8% African American, 7.8% Asian, and 11.8% other. Each participant was asked to respond to 3-4 different representative heuristic questions, providing three responses for each ReHeAT question. We used majority voting to finalize the human decision. It can be observed that the human results closely matched the LLaMA2-70B result, while the GPT series performed slightly better than the human results. This aligns with the findings of Bar-Hillel & Neter (1993), Kahneman & Tversky (1973), Kahneman et al. (1982), Tversky & Kahneman (1974), and Tversky & Kahneman (1983), who reported that human accuracy across various tasks ranges from 10% to 50%.
Type Model
Human GPT-3.5 GPT-4 LLaMA2-70B PaLM 2
Base Rate Fallacy 27.7 31.9 (+4.2) 34.0 (+6.3) 23.4 (-4.3) 19.1 (-8.6)
Conjunction Fallacy 24.4 31.1 (+6.7) 20.0 (-4.4) 20.0 (-4.4) 2.2 (-22.2)
Disjunction Fallacy 12.5 22.9 (+10.4) 6.3 (-6.2) 10.4 (-2.1) 2.3 (-10.2)
Insensitivity to Sample Size 33.3 45.1 (+11.8) 54.9 (+21.6) 43.1 (+9.8) 22.0 (-11.3)
Misconceptions of Chance 50.0 50.0 (=) 50.0 (=) 75.0 (+25.0) 50.0 (=)
Regression Fallacy 85.7 71.4 (-14.3) 71.4 (-14.3) 71.4 (-14.3) 71.4 (-14.3)
Average 27.2 34.7 (+7.5) 31.2 (+4.0) 27.2 (=) 14.8 (-12.4)
Table 3: Human performance on ReHeAT. A comparison with standard results from LMs.
Overall, LLMs have demonstrated a representativeness heuristic bias similar to humans. Example questions and model answers are in Table 18 - 23 of the Appendix.
@@ 5.2 Discrepancies in Model Reasoning: Beyond Predictions
Generally, GPT-4 outperformed LLaMA2-70B in all types of RHQs. However, this improvement is accompanied by a side effect: an increased proportion of instances where GPT-4’s reasoning process is accurate, but the outcome is incorrect. This primarily occurs because when faced with ambiguous questions, GPT-4 often preferred to express the need for additional information before making a decision. This is also why GPT-4 accuracy is slightly lower than GPT-3.5 under some prompts. Although we mark such cautious answers as incorrect outcomes—because they did not adhere to the directive of providing a definitive answer—cautious answers may remain beneficial in practical settings. Conversely, LLaMA2-70B more frequently produces correct outcomes from incorrect reasoning processes; the reasons are often incorrectly based on a stereotype rather than the problem’s description.
We have conducted case studies on some interesting reasoning, we include the case study in Appendix A. In summary, our findings indicate that LLMs are influenced by semantic content in their reasoning processes, aligning with results from previous research (Acerbi & Stubbersfield, 2023; Dasgupta et al., 2022). Specifically, when the semantic content of tasks was clear and did not contain sensitive content (e.g. gender, race), the models’ responses were better.
@@
5.3 Improving Performance by Hinting LLMs to Use Their Knowledge
How can the model’s cognitive process be put back on track? We test whether LLMs show enhanced performance with prompts that hint for them to use their existing knowledge. We tested two types of hints: one general and the other based on more detailed cues for each representativeness heuristic type (Table 17 of the Appendix). These prompts aimed to hint at the model to recall the knowledge it possessed. Results are presented in Table 4. We also provide detailed results for each type of representativeness heuristic in Table 12 and Table 13 of the Appendix. We found that both types of hints provided a noticeable improvement for most models, with the specific type of hints providing a more significant boost in performance compared to the general prompts.
@@ 5.4 The Impact of Situational Similarity on Model Performance
After observing the performance differences between the model on statistical SPQs and on the contextual RHQ counterparts found in ReHeAT, we delve further into the middle of these two questions - Intermediate Questions (IQs) (Table 16 of the Appendix). This type of question integrates specific situational contexts and statistical data, making it necessary to consider both concrete data and the potential impact of the situation when making decisions. For example, for Linda’s problem, we assign a probability to each independent event to offer the model with more explicit statistical information. Nevertheless, the model has to infer the information based on the similarity manifested in the meaning of the statements in RHQs. The performance of the models on IQs is reported in Table 6 of the Appendix. Compared with SPQs’ results (Table 5 of the Appendix), more errors were made for IQs. This further indicates that introducing scenarios interferes with the model’s statistical decision-making process. This also explains why LLMs exhibit the representativeness heuristic more frequently when responding to RHQs.